OpenStudy (anonymous):

simplify : (ln2/(24110*364.25*24*60*60))

OpenStudy (anonymous):

ln2 = 0.69314718055994530941723212145818 (24110*364.25*24*60*60)) = 758770632000 (ln2/(24110*364.25*24*60*60)) = 0.69314718055994530941723212145818/ 758770632000 = 9.1351345364134402795024375885199e-13